OpenStudy (1dfanatic):

Solve by elimination. 2x + y = 3 -2x + y = 1

OpenStudy (photon336):

Well, elimination we can combine both equations by adding them

OpenStudy (photon336):

This means we must add everything together, but generally speaking sometimes it's not that straightforward. in our example we can eliminate one variable so that we can solve for the other variable. so what would be the result of this? you'll see the "elimination" part one variable vanishes. tell me what you get for this -2x+2x+y+y = 3+1

OpenStudy (1dfanatic):

2?

OpenStudy (photon336):

how did you get that?

OpenStudy (1dfanatic):

well the 2's cancel out and then y+y=2 and 3+1=4 then you divide and 4/2 is 2

OpenStudy (photon336):

:) yep so y = 2 but you're not done yet

OpenStudy (photon336):

Now that you know that y = 2, you need to find what x is. have any ideas of how to do this?

OpenStudy (1dfanatic):

i think subtract the whole first equation from the first??

OpenStudy (photon336):

so @1Dfanatic we know now that y = 2 right? what we can do now is easy. wherever we see y we plug in 2 and then we can solve for x. 2x+y = 3 make sense?

OpenStudy (1dfanatic):

4+y=3 right?

OpenStudy (photon336):

no, we know Y = 2 so we plug in y = 2 wherever we see y in our equation

OpenStudy (1dfanatic):

so 2+2=3?

OpenStudy (photon336):

Take for example 2x+y = 3, y = 2

OpenStudy (photon336):

I plugged in 2 where y was because we found before that y = 2 So it becomes 2x+2 = 3

OpenStudy (photon336):

@1Dfanatic

OpenStudy (1dfanatic):

oooh then it'd be 1/2

OpenStudy (photon336):

x = 1/2 and y = 2

OpenStudy (photon336):

now you can actually check to see if these answers are correct and if they are, they should satisfy BOTH equations

OpenStudy (photon336):

Try plugging in both x =1/2 and y = 2 into the bottom equation too -2x+y = 1 tell me if it's true

OpenStudy (1dfanatic):

Either i didn't do my math right or it;s not right,,